How do you solve:
2^3x-2 = √128

Answer by algebrahouse.com(1659)

2^(3x - 2) = √128
2^(3x - 2)= 2^(7/2) {changed 128 to have a base of 2 and an exponent of 7/2}
3x - 2 = 7/2 {since bases are equal, set the exponents equal to each other}
3x = 11/2 {added 2 to each side}
x = 11/6 {divided each side by 3}

For more help from me, visit: www.algebrahouse.com